OpenStudy (amtran_bus):

How do you solve a multistep trig equation? Sin (3 Theta) =-1

OpenStudy (amtran_bus):



OpenStudy (ghazi):

\[\sin 3 \theta =-1\] now taking inverse of it \[3 \theta = \sin^{-1}(-1)\] and sin is -1 at 3 pi/2

OpenStudy (amtran_bus):

Right. That is what I am getting also. But what do I need to do after that? My book is giving four crazy answers.

OpenStudy (ghazi):

divide both sides by 3

OpenStudy (amtran_bus):